OpenStudy (kewlgeek555):

I Need Someone Good with Linear Functions, please, to check my work and solve some questions.

OpenStudy (kewlgeek555):

Please post to let me know.

OpenStudy (kewlgeek555):

Anyone willing to help me? Please comment...

OpenStudy (anonymous):

yes, bro, it's a help site :) post it

OpenStudy (kewlgeek555):

Okay. Thanks. So did I do this right? Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1) I chose B

OpenStudy (anonymous):

\[(y - y_1) = m(x - x_1)\] m is the slope and (x1, y1) represents any point on the line the answer would be C since the slope is one half

OpenStudy (kewlgeek555):

Okay. Thanks. I have another one I want you to check please. Sorry, this is a big test, but at least I am trying! The graph below shows four straight lines, P, Q, R, and S. http://orange.flvs.net/webdav/assessment_images/educator_algebra1_v17/03_00_pretest/03_00_32.jpg Which line is represented by the function f(x) = 4x - 3? Line P Line Q Line R Line S I chose D, line S because it said minus 3, which would also be -3 in the y-intercept.

OpenStudy (kewlgeek555):

@Euler271

OpenStudy (kewlgeek555):

@agent0smith

OpenStudy (kewlgeek555):

@mathstudent55

OpenStudy (kewlgeek555):

@mathslover

OpenStudy (kewlgeek555):

@mathslover please ignore hte conversation - I still have that question.

mathslover (mathslover):

Please post it as another question. Someone will come and help you out. Cheers

mathslover (mathslover):

Well but yes, you're right it is D)

OpenStudy (kewlgeek555):

Thanks. Do you think you could stick around and help? Or do you have ambassador duties?

mathslover (mathslover):

Well we do both, being an ambassador is a voluntary work, we don't get any pay for it. We are just to help new users with OpenStudy's working and some features. We invite more and more users to this site so that more students can take advantage of this site. We don't have any time consuming ambassador duties, we can stick around and help also.

OpenStudy (kewlgeek555):

Oh, okay. Well I wanted to know about this question: Choose the equation of the horizontal line that passes through the point (-8, -7). y = -8 y = -7 x = -8 x = -7

OpenStudy (agent0smith):

vertical line means x=... horizontal line means y=... so you can already eliminate two choices. The line has to pass through (-8, -7) (ie has to pass through -7 on the y-axis) so which of the remaining two choices will passes through -7 on the y axis?